### v2

@@ III Proposed PAPR Reduction Scheme
During each iteration, the pre-chirp values within a designated group undergo complete reversal. Then, a new $\mathrm{PAPR}^{(l)}$ is calculated using an updated pre-chirp matrix $\mathbf{\Lambda}_{c2}^{(l)}$. Upon completion of the algorithm, the minimum PAPR among $(V+1)$ values, denoted as $\mathrm{PAPR}_{\min}$, along with its corresponding pre-chirp matrix $\mathbf{\Lambda}_{c_{2}}^{*}$, is selected as the output.
Algorithm 1 GPS with $W=2$ in AFDM System for PAPR Reduction
0:  Number of subcarriers $N$, number of groups $V$, number of subcarriers per group $M$, post-chirp matrix $\mathbf{\Lambda}_{c_{1}}$, normalized FFT matrix $\mathbf{F}$, symbol vector ${\mathbf{x}}$;
2:  $\mathrm{PAPR}_{\min}=\mathrm{PAPR}^{(0)}$;
3:  for $l=1:V$ do
4:     $\mathbf{\Lambda}_{c_{2}}^{(l)}=\mathbf{\Lambda}_{c_{2}}^{(l-1)}$;
5:     $\mathbf{\Lambda}_{c_{2}}^{(l)}\left((l\!-\!1)M\!+\!1\!:\!lM\right)\!=\!\mathbf{\Lambda}_{c_{2}}^{(l)}\left((l\!-\!1)M\!+\!1\!:\!lM\right)\!^{H}$;
6:     Calculate $\mathrm{PAPR}^{(l)}$ similar to (24);
7:     if $\mathrm{PAPR}^{(l)}\geq\mathrm{PAPR}_{\min}$ then
8:        $\mathbf{\Lambda}_{c_{2}}^{(l)}\!\left((l\!-\!1)M\!+\!1\!:\!lM\right)\!\!=\!\!\mathbf{\Lambda}_{c_{2}}^{(l)}\!\left((l\!-\!1)M\!+\!1\!:\!lM\right)\!^{H}$;
9:     else
10:        $\mathrm{PAPR}_{\min}=\mathrm{PAPR}^{(l)}$;
11:     end if
12:  end for
13:  $\mathbf{\Lambda}_{c_{2}}^{*}=\mathbf{\Lambda}_{c_{2}}^{(V)}$;
13:  $\mathrm{PAPR}_{\min}$ and $\mathbf{\Lambda}_{c_{2}}^{*}$.
@@ IV Numerical Results
An AFDM system with 16QAM under LTV channel is considered. Without loss of generality, channel coefficients $h_{i}\sim\mathcal{CN}\left(0,1\right)$ are generated. The number of subcarriers, elements per set, and groups are set as $N=64$, $W=2,3,4$, and $V=4,8,16$, respectively.
Figure 3: PAPR performance of 16QAM AFDM system with GPS for different $W$ and $\Omega_{m}$.
@@
In Fig. 3, simulation results of GPS are presented for different values and number of elements in $\Omega_{m}$, where $V=4$ and adjacent grouping pattern are applied. In terms of values of elements, the results demonstrate that a larger phase difference yields improved performance, assuming an equivalent number of elements. In terms of number of elements $W$, the performance gain from the increase of number converges quickly. Compared with the original AFDM, a reduction in PAPR by approximately $2.5$ dB can be achieved when $W=2$ and $\Omega_{m}=[\frac{1}{4m^{2}},-\frac{1}{4m^{2}}]$.
Figure 4: PAPR performance of 16QAM AFDM system with GPS for different $V$ and grouping patterns.
@@
Fig. 4 illustrates the impact of the number of groups and different grouping strategies on performance, where $W=2$ and $\Omega_{m}=[\frac{1}{4m^{2}},-\frac{1}{4m^{2}}]$ are employed. It can be observed that adjacent grouping consistently outperforms comb grouping. Targeting the CCDF at a level of $10^{-4}$, the adjacent grouping pattern demonstrates performance improvements of approximately $0.7$ dB, $0.3$ dB, and $0.1$ dB over the comb grouping pattern for configurations with $4$, $8$, $16$ groups, respectively. In addition, increasing the number of groups could enhance the effectiveness of PAPR reduction. However, the rate of performance improvement diminishes as the number of groups increases, which indicates a significant increase in complexity. To balance complexity and performance, it is advisable to keep the number of groups minimal, provided that the desired PAPR reduction threshold is achieved.
@@
Figure 5: PAPR performance of 16QAM AFDM system with GPS and enumerated algorithm.
@@
In Fig. 5, the performance disparity between the enumerated algorithm and the GPS is depicted, where $W=2$, $\Omega_{m}=[\frac{1}{4m^{2}},-\frac{1}{4m^{2}}]$, and the adjacent grouping pattern are applied. The enumerated algorithm consistently outperforms the non-enumerated algorithm. However, this enhancement comes at the cost of significantly increased computational complexity. Specifically, the computational complexity is $\frac{2^{V}}{V}$ times higher in the case of $V$ groups. Notably, the GPS with $V=8$ achieves a performance improvement of approximately $0.4$ dB over the enumerated algorithm with $V=4$, while having lower computational complexity. Consequently, the non-enumerated GPS is deemed more efficient in practical implementations.
@@ V Conclusion
In this letter, a GPS algorithm is proposed to reduce PAPR in AFDM. We first modify the pre-chirp module of conventional AFDM, ensuring that such an adjustment preserves the inherent characteristics of AFDM. Subsequently, we discuss several factors in GPS to achieve the optimal balance between performance and complexity. Finally, simulations are conducted to verify the performance of GPS. The numerical results show that the adjacent grouping pattern exhibits superior performance compared to the comb grouping pattern. Other better grouping patterns and the allocation of the number of subcarriers per group could be interesting directions for future research.

### v4

@@ III-3 Grouping
During each iteration, the pre-chirp values within a designated group undergo complete reversal. Then, a new $\mathrm{PAPR}^{(l)}$ is calculated using an updated pre-chirp matrix $\mathbf{\Lambda}_{c2}^{(l)}$. Upon completion of the algorithm, the minimum PAPR among $(V+1)$ values, denoted as $\mathrm{PAPR}_{\min}$, along with its corresponding pre-chirp matrix $\mathbf{\Lambda}_{c_{2}}^{*}$, is obtained as the output.
@@ IV Numerical Results
An AFDM system with 16QAM under LTV channel is considered. Without loss of generality, channel coefficients $h_{i}\sim\mathcal{CN}\left(0,1\right)$ are generated. The number of subcarriers is set as $N=64$. At the receiver, the channel information is assumed to be perfectly known. Totally, $\lceil{\log_{2}(W^{V})}\rceil$ extra side bits are needed for decoding, and the minimum mean square error (MMSE) equalization is applied.
(a) Different $W$ and $\Omega_{m}$
(b) Different $V$ and grouping patterns
Figure 3: PAPR performance of 16QAM AFDM with GPS.
@@
In Fig. 3a, simulation results of GPS are presented for different values and number of elements in $\Omega_{m}$, where $V=4$ and adjacent grouping pattern are applied. In terms of values of elements, the results demonstrate that a larger phase difference yields improved performance, assuming an equivalent number of elements, which aligns with the conclusion derived from our derivation. In terms of number of elements $W$, the performance gain from the increase of number converges quickly.
@@
Fig. 3b illustrates the impact of the number of groups and different grouping strategies on performance, where $W=2$ and $\Omega_{m}=[\frac{1}{4m^{2}},-\frac{1}{4m^{2}}]$ are employed. It can be observed that adjacent grouping consistently outperforms comb grouping. Targeting the CCDF at a level of $10^{-4}$, the adjacent grouping pattern demonstrates performance improvements of approximately $0.7$ dB, $0.3$ dB, and $0.1$ dB over the comb grouping pattern for configurations with $4$, $8$, $16$ groups, respectively. In addition, increasing the number of groups could enhance the effectiveness of PAPR reduction. However, the rate of performance improvement diminishes as the number of groups increases, which indicates a significant increase in complexity. To balance complexity and performance, it is advisable to keep the number of groups small, provided that the desired PAPR reduction threshold is achieved.
@@
Figure 4: PAPR performance of 16QAM AFDM with GPS, enumerated algorithm, and OTFS.
@@
In Fig. 4, the performance disparity between the enumerated method and the GPS is depicted, where $W=2$, $\Omega_{m}=[\frac{1}{4m^{2}},-\frac{1}{4m^{2}}]$, and the adjacent grouping pattern are applied. The enumerated method consistently outperforms the GPS. However, this enhancement comes at the cost of significantly increased computational complexity. Notably, the GPS with $V=8$ ($\mathcal{O}\left(8\right)$) achieves a performance improvement of approximately $0.4$ dB over the enumerated algorithm with $V=4$ ($\mathcal{O}\left(16\right)$), while having lower computational complexity. Consequently, the non-enumerated GPS is deemed more efficient in practical implementations. Additionally, a comparison with OTFS shows that AFDM with GPS ($V=2$) achieves better PAPR performance while having lower modulation complexity, since OTFS employs two-dimensional modulation whose complexity is significantly higher than the one-dimensional modulation in AFDM.
Figure 5: BER performance of 16QAM AFDM with GPS.
Fig. 5 presents the bit error rate (BER) performance of AFDM with GPS achieved with additional side information, which is found to be closely comparable to the original AFDM. Targeting a BER performance of $10^{-5}$, the signal-to-noise ratio (SNR) loss resulting from the GPS is about $0.25$ dB. With prior knowledge of the side information, there is no BER performance degradation, confirming that AFDM with GPS preserves the properties of the original AFDM. Additionally, when $N=64$, $V=4$, and $W=2$, the spectral efficiency is about $98.5\%$ of that in original AFDM. To sum up, the results indicate that the GPS exerts a slight impact on the communication performance.
@@ V Conclusion
In this letter, a GPS algorithm is proposed to reduce PAPR in AFDM. We first modify the pre-chirp module of conventional AFDM, ensuring that such an adjustment preserves the inherent characteristics of AFDM. Subsequently, we discuss three aspects in PAPR reduction scheme to achieve the optimal balance between performance and complexity. Finally, simulations are conducted to verify the effectiveness of the proposed GPS. With an acceptable increase in complexity and a marginal degradation in communication performance, AFDM with GPS can achieve a lower PAPR than OTFS, while maintaining a lower modulation complexity.